Money is invested in a savings account at 4% simple interest. After 1 year, there is $364 in the account. How much was originally invested?
I can never remember the formula for this type of equation. I know you have to either multiply or divide the % by the $364 to get the amount of interest per month, and then multiply that number by 12 for the 12 months. And then subtract that number from the $364 to find the original amount. I just don't know that first step.

Answer by addingup(3677)

Future Value= Present Value(1+rt)where:
r= interest rate
t= number of years
364= PV(1+0.04(1))=
364= PV(1+0.04)=
364= PV(1.04) Divide both sides by 1.04:
PV= 364/1.04=
PV= 350 is how much was originally invested.